Daily Archives: November 14, 2012

Spectrometers Back Up

The problem of the building air control system has been fixed. All the NMR spectrometers are up running at 11:30pm. Thank Sarah Lindahl for timely reporting the problem.

Posted in Announcements | Comments Off

Emergency Spectrometer Shutdown

The room (C237) is getting very warm (> 40 C) duo to a problem on the building air control system. All the spectrometers in the chemistry NMR lab have been powered off at 8pm.

Posted in Announcements | Comments Off